# Hyper compressed fine-tuning of large foundation models with Quantum Inspired Adapters

### v1

@@ Abstract
Fine-tuning pre-trained large foundation models for specific tasks has become increasingly challenging due to the computational and storage demands associated with full parameter updates. Parameter-Efficient Fine-Tuning (PEFT) methods address this issue by updating only a small subset of model parameters using adapter modules. In this work, we propose Quantum-Inspired Adapters, a PEFT approach inspired by Hamming-weight preserving quantum circuits from quantum machine learning literature. These models can be both expressive and parameter-efficient by operating in a combinatorially large space while simultaneously preserving orthogonality in weight parameters. We test our proposed adapters by adapting large language models and large vision transformers on benchmark datasets. Our method can achieve 99.2% of the performance of existing fine-tuning methods such LoRA with a 44x parameter compression on language understanding datasets like GLUE and VTAB. Compared to existing orthogonal fine-tuning methods such as OFT or BOFT, we achieve 98% relative performance with 25x fewer parameters. This demonstrates competitive performance paired with a significant reduction in trainable parameters. Through ablation studies, we determine that combining multiple Hamming-weight orders with orthogonality and matrix compounding are essential for performant fine-tuning. Our findings suggest that Quantum-Inspired Adapters offer a promising direction for efficient adaptation of language and vision models in resource-constrained environments.
@@ 1 Introduction
Pre-trained large foundation models such as BERT [1], GPT-3 [2], and Vision Transformers [3] have achieved state-of-the-art results on various tasks. Fine-tuning these models on specific downstream tasks typically involves updating all model parameters but with a lower learning rate, which becomes computationally prohibitive as model sizes continue to grow into the billions of parameters. This challenge has spurred interest in Parameter-Efficient Fine-Tuning (PEFT) methods [4], which aim to adapt large foundation models to new tasks by updating only a small subset of parameters or introducing lightweight adaptation modules.
@@
One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) [5], which injects low-rank trainable matrices into the transformer layers, significantly reducing the number of parameters that need to be updated. Other methods like Adapters [4], BitFit [6], and Prompt Tuning [7] have also demonstrated effectiveness in various settings. Recently, Orthogonal Fine-Tuning (OFT) [8] and its ‘Butterfly’ specification (BOFT) [9] have been proposed to mitigate catastrophic forgetting of the pre-trained models during fine-tuning by applying orthogonal transformations. These methods have shown promising results in achieving a balance between performance and parameter efficiency.
@@
Figure 1: Comparison of different adapter methods. Trainable parameters for each model shown in dark green. a) Full finetuning, i.e. retraining the entire pre-trained weight matrix. b) Low-rank adaptation (LoRA) using an additive adapter decomposition into two low-rank matrices. c) Orthogonal fine-tuning (OFT) with a multiplicative adapter decomposed into orthogonal blocks. d) Our proposed specification of a (multiplicative) Quantum-Inspired Adapter, using compound matrices decomposed (like OFT) into blocks. For the compound matrices, the zeroth order compound (top left of each block) is the only trainable part. Higher order compounds are completely determined by this base matrix.
@@
In this work, we introduce Quantum-Inspired Adapters, a novel PEFT method inspired by Hamming-weight preserving quantum circuits [10, 11, 12]. Our approach constructs orthogonal adapters using compound matrices, focusing on compound orders up to a certain value $k$ to ensure parameter efficiency. We evaluate our method on several tasks from the general language understanding evaluation (GLUE) benchmark  [13] and on a subset of tasks from the visual task adaptation (VTAB) benchmark [14]. Our experiments demonstrate that Quantum-Inspired Adapters achieve competitive performance while dramatically reducing the number of trainable parameters compared to existing PEFT methods like LoRA, OFT, and BOFT.
Our contributions are summarized as follows:
  • We propose Quantum-Inspired Adapters, a novel PEFT method that leverages circuits from quantum machine learning literature to create parameter efficient adapters.
  • We provide a detailed classically-efficient construction of quantum inspired adapters using compound matrices and showcase parameter efficiency and orthogonality properties.
  • We conduct extensive experiments on the GLUE and VTAB benchmark, demonstrating that our method achieves a significant reduction in trainable parameters while maintaining competitive performance levels.
  • We analyze the trade-offs between accuracy and parameter efficiency and discuss the implications for model adaptation in resource-constrained environments.

### v2

# QuIC: Quantum-Inspired Compound Adapters for Parameter Efficient Fine-Tuning

@@ Abstract
Scaling full finetuning of large foundation models strains GPU memory and training time. Parameter Efficient Fine-Tuning (PEFT) methods address this issue via adapter modules which update only a small subset of model parameters. In this work, we introduce Quantum-Inspired Compound Adapters (QuIC Adapters), a PEFT approach inspired from Hamming-weight preserving quantum circuits that can effectively finetune a model using less than $0.02\%$ memory footprint of the base model. QuIC adapters preserve pretrained representations by enforcing orthogonality in weight parameters, and have native deployment mechanisms on quantum computers. We test QuIC adapters by finetuning large language models like LLaMA and vision transformers on language, math, reasoning and vision benchmarks. In its first-order configuration, QuIC recovers the performance of existing orthogonal methods, while higher-order configurations enable substantial parameter compression (over $40\times$ smaller than LoRA) for a modest performance trade-off, unlocking applications in highly resource-constrained environments. Through ablation studies, we determine that combining multiple Hamming-weight orders with orthogonality and matrix compounding are essential for performant finetuning. Our findings suggest that QuIC adapters offers a promising direction for efficient finetuning of foundation models in resource-constrained environments.
@@ 1 Introduction
Pre-trained large foundation models such as BERT [Devlin et al., 2018], GPT-3 [et al., 2020], and Vision Transformers [Dosovitskiy, 2020] have achieved state-of-the-art results on various tasks. Finetuning these models on specific downstream tasks typically involves updating all model parameters but with a lower learning rate, which becomes computationally prohibitive as model sizes continue to grow into the billions of parameters. This challenge has spurred interest in Parameter-Efficient Fine-Tuning (PEFT) methods [Houlsby et al., 2019], which aim to adapt large foundation models to new tasks by updating only a small subset of parameters or introducing lightweight adaptation modules.
@@
One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) [Hu et al., 2021], which injects low-rank trainable matrices into transformer layers, significantly reducing the number of parameters that need to be updated. Other methods like Adapters [Houlsby et al., 2019], BitFit [Ben Zaken et al., 2022], and Prompt Tuning [Lester et al., 2021] have also demonstrated effectiveness in various settings. Recently, Orthogonal Fine-Tuning (OFT) [Qiu et al., 2023] and its ‘Butterfly’ specification (BOFT) [Liu et al., 2023] have been proposed to mitigate catastrophic forgetting of the pre-trained models during finetuning by applying orthogonal transformations. These methods have shown promising results in achieving a balance between performance and parameter efficiency.
While methods like LoRA and OFT significantly reduce parameters compared to full finetuning, a critical need remains for even greater efficiency in resource-constrained scenarios. Deploying personalized models on-device (e.g., smartphones or wearables), serving thousands of task-specific adapters simultaneously, or reducing bandwidth in federated learning all impose strict memory and storage budgets that even conventional PEFT methods can exceed [Kopiczko et al., 2024, Zhang et al., 2023, YEH et al., 2024]. This motivates the development of methods capable of extreme compression, pushing the Pareto frontier of what is achievable with a minimal parameter budget.
@@
Figure 1: Comparison of different adapter methods. Trainable parameters for each model shown in dark green. a) Full finetuning b) Low-rank adaptation (LoRA) c) Orthogonal finetuning (OFT) d) Quantum-Inspired Compound adapter (QuIC adapter). For QuIC adapters, the zeroth order compound (top left of each block) is the only trainable part. Higher order compounds are completely determined by this base matrix.
Inspired by the potential exponential compression abilities of quantum and quantum-inspired computing, there has been a growing interest in Quantum-Inspired PEFT methods such as QuanTA [Chen et al., 2024] and QPA [Liu et al., 2025]. While QuanTA constructs adapters via contracted quantum-inspired tensor networks, Quantum Parameter Adaptation (QPA) uses quantum circuits to generate parameters for methods such as LoRA. These works highlight the potential for quantum machine learning within finetuning, however both methods contain a number of bottlenecks which potentially prohibit quantum computer integration with finetuning pipelines at larger scales.
@@
In this work, we propose Quantum-Inspired Compound Adapters (QuIC Adapters), a novel PEFT method inspired by Hamming-weight preserving quantum circuits [Kerenidis and Prakash, 2022, Landman et al., 2022, Cherrat et al., 2023]. With QuIC adapters, orthogonality is a native feature, and we focus on compound orders up to a certain constant $K$ to ensure parameter efficiency. We evaluate our method on several datasets over a variety of domains. For language, vision, reasoning and math problems, we use the the General Language Understanding Evaluation (GLUE) benchmark [Wang, 2018], a subset of tasks from the Visual Task Adaptation (VTAB) benchmark [Zhai et al., 2019], the Discrete Reasoning Over the text in the Paragraph (DROP) dataset [Dua et al., 2019], and the MATH10K [Hu et al., 2023] benchmark respectively. On the model side, we finetune the moderate size DeBERTaV3 [He et al., 2021] for language and DINOv2-large for vision. For a larger model and for math and reasoning tasks we focus on LLaMA-7B [Touvron et al., 2023b]. Our experiments demonstrate that QuIC adapters achieve competitive performance while dramatically reducing the number of trainable parameters compared to existing PEFT methods like LoRA, OFT, BOFT and QuanTA, among others.